OpenStudy (anonymous):

Write the equation of a line that passes through the point (-2, 1) and has a slope of 4.

pooja195 (pooja195):

In what form?

OpenStudy (anonymous):

y=mx+b

OpenStudy (anonymous):

i know y=4x

pooja195 (pooja195):

ok well we need to find the y intercept :) but lets just do everything at once :P see this formula? Plug in the given info and then solve for y :) \[\huge y-y_1=m(x-x_1)\] where m=slope (x1,y1)

OpenStudy (anonymous):

y=4x+1?

pooja195 (pooja195):

No try again can you show your work? So i can see where you made the mistake?